OpenStudy (anonymous):

Solve the system of equations using matrices. Use Gauss-Jordan elimination 3x-7y-7z = 7 6x+4y-3z=67 -6x-3y+z=-62

OpenStudy (across):

Hint:\[\begin{bmatrix} 3 & -7 & -7\\ 6 & 4 & -3\\ -6 & -3 & 1 \end{bmatrix}\begin{bmatrix} x\\ y\\ z \end{bmatrix}=\begin{bmatrix} 7\\ 67\\ -62 \end{bmatrix}.\]

OpenStudy (anonymous):

so does x = 7

OpenStudy (nincompoop):

hmm

OpenStudy (nincompoop):

did you use Michael Jordan Elimination method to arrive at x=7?

OpenStudy (anonymous):

the book I am using is really confusing it says I have to get 1 on top but I don't know how

OpenStudy (nincompoop):

http://ceee.rice.edu/Books/CS/chapter2/linear44.html

OpenStudy (anonymous):

I see how to do it when the number is 0 or 1 but I don't know how to start with the 3

OpenStudy (anonymous):

@across almost gave you the cramers rule use it to find values of x, y, and z

OpenStudy (nincompoop):

http://www.karlscalculus.org/cgi-bin/linear.pl

OpenStudy (anonymous):

okay I used that but my answers I think are what is confusing me these are my choices a. {(7,1,7)} b. {14,7,-7)} c. {(-7,7,14)} d. {(7,7,1)}

OpenStudy (across):

If you have choices, all you have to do is substitute, I mean, slam-dunk them and check!

OpenStudy (anonymous):

I did that and none of them equal up

OpenStudy (anonymous):

that is the way they are

hero (hero):

I'm interested in knowing that this Michael Jordan method is